# Stationary Policies are Optimal in Risk-averse Total-reward MDPs with EVaR

### v1

@@ 3 Solving ERM Total Reward Criterion
Our objective in this section is to maximize the ERM of the infinite-horizon total sum of rewards for some given $\beta>0$:
$\sup_{\pi\in\Pi_{\mathrm{HR}}}\liminf_{t\to\infty}\operatorname{ERM}_{\beta}^{\pi,\bm{\mu}}\left[\sum_{k=0}^{t}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right].$ (5)
Analogously to (1), the superscripts indicate the policy and initial state distribution that govern $\tilde{s}_{t}$ and $\tilde{a}_{t}$. The definition employs limit inferior because the limit may not exist for non-stationary policies.
To streamline the notation, we define $g_{t}\colon\Pi_{\mathrm{HR}}\times\mathbb{R}_{++}\to\mathbb{R}$ and $g_{t}^{\star}\colon\mathbb{R}_{++}\to\bar{\mathbb{R}}$ as
$g_{t}(\pi,\beta)$ $:=\operatorname{ERM}_{\beta}^{\pi,\bm{\mu}}\left[\sum_{k=0}^{t}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]$ (6)
$g_{t}^{\star}(\beta)$ $:=\sup_{\pi\in\Pi_{\mathrm{HR}}}g_{t}(\pi,\beta).$
We also define $g_{\infty}(\pi,\beta):=\liminf_{t\to\infty}g_{t}(\pi,\beta)$ and $g_{\infty}^{\star}(\pi,\beta):=\liminf_{t\to\infty}g_{t}^{\star}(\beta)$. Note that the functions $g_{\infty}$ and $g_{\infty}^{\star}$ can return infinite values and that (5) differs from $g^{\star}_{\infty}$ in the order of the limit and the supremum. Finally, for the border case when $\beta=0$, we assume that all $g$ functions are defined in terms of the expectation operator.
In the remainder of the section, we assume that the risk level $\beta>0$ is fixed and omit it in notations when its value is unambiguous from the context.
@@ 3.1 Finite Horizon
Finite-horizon functions $\bm{v}^{t}(\pi)\in\mathbb{R}^{S}$ and $\bm{v}^{t,\star}\in\mathbb{R}^{S}$ are defined for each horizon $t\in\mathbb{N}$ and policy $\pi\in\Pi_{\mathrm{MD}}$, $s\in\bar{\mathcal{S}}$ as
@@
$v^{t,\star}_{s}$ $\;:=\;\max_{\pi\in\Pi_{\mathrm{MD}}}v^{t}_{s}(\pi).$
Because the nonlinearity of ERM complicates the analysis, it will be convenient to instead rely on exponential value function $\bm{w}^{\pi}\in\mathbb{R}^{S}$ for $\pi\in\Pi_{\mathrm{MD}}$, $t\in\mathbb{N}$, and $s\in\mathcal{S}$ that satisfy
@@
The optimal $\bm{w}^{t,\star}\in\mathbb{R}^{S}$ is defined analogously from $\bm{v}^{t,\star}$. Note that $\bm{w}^{t}<\bm{0}$ (componentwise) and $\bm{w}^{0}(\pi)=\bm{w}^{t,\star}=-\bm{1}$ for any $\pi\in\Pi_{\mathrm{MD}}$. Similar exponential value functions have been used previously in exponential utility function objectives (Denardo and Rothblum 1979; Patek 2001) and in the analysis of robust MDPs (Grand-Clément and Petrik 2022).
@@
$L^{\star}\bm{w}$ $\;:=\;\max_{\bm{d}\in\mathcal{D}}L^{\bm{d}}\bm{w}=\max_{\bm{d}\in\operatorname{ext}\mathcal{D}}L^{\bm{d}},$
@@
For each $t=1,\dots$, and $\pi=(\bm{d}_{0},\dots,\bm{d}_{t-1})\in\Pi_{\mathrm{MR}}^{t}$ the exponential values satisfy that
@@
where $\exists\pi^{\star}\in\Pi_{\mathrm{MD}}^{t}$.
The proof of Theorem 3.1 result is standard and has been established both in the context of ERMs (Hau, Petrik, and Ghavamzadeh 2023) and utility functions (Patek 1997).
The following important result follows directly from Theorem 3.1 by algebraic manipulation and by the monotonicity of exponential value function transformation and the ERM.
@@ 3.2 Infinite Horizon
We now turn to constructing infinite-horizon optimal policies as a limiting case of the finite horizon. An important quantity is the infinite-horizon exponential value function defined for each $\pi\in\Pi_{\mathrm{HR}}$ as
$\bm{w}^{\infty}(\pi)\;:=\;\liminf_{t\to\infty}\bm{w}^{t}(\pi),\quad\bm{w}^{\infty,\star}(\pi)\;:=\;\liminf_{t\to\infty}\bm{w}^{t,\star}.$
Note again that we use the inferior limit because the limit may not be defined for non-stationary policies.
The following theorem is the main result of this section. It shows that in the infinite horizon, the optimal exponential value function is attained by a stationary deterministic policy and is a fixed point of the exponential Bellman operator.
@@
We get the following immediate corollary of Theorem 3.3 that characterizes the exponential value functions.
@@
There exists a $\pi^{\star}\in\Pi_{\mathrm{SD}}$, such that $\bm{v}^{\infty,\star}=\bm{v}^{\infty}(\pi^{\star})$ and
@@
We now outline the proof of Theorem 3.3; please see Section D.4 for details. To establish Theorem 3.3, we show that $\bm{w}^{t,\star}$ converges to a fixed point as $t\to\infty$. Standard arguments do not apply to our setting (Puterman 2005; Kallenberg 2021; Patek 2001) because the ERM-TRC Bellman operator is not an $L_{\infty}$-contraction, is not linear, and the values in value iteration do not increase or decrease monotonically. Although the exponential Bellman operator $L^{\bm{d}}$ is linear, it may not be a contraction either.
@@
The value $\bm{z}$ can be interpreted as the exponential value function at the termination of the process following $\pi$ for $t$ periods, with exponential value at termination $\bm{z}$. Note that $\bm{w}^{t}(\pi)=\bm{w}^{t}(\pi,\bm{1}),\,\forall\pi\in\Pi_{\mathrm{MR}},t\in\mathbb{N}.$
An important technical result we show is that the only way a stationary policy’s return can be bounded is if the policy’s matrix has a contracting spectral radius.
@@
Lemma 3.5 uses the transience property to show that Perron vector (with the maximum absolute eigenvalue) $\bm{f}$ of $\bm{B}^{\bm{d}}$ satisfies that $\bm{f}^{\top}\bm{b}^{\bm{d}}>0$. Therefore, $\rho(\bm{B}^{\bm{d}})<1$ is necessary for the series in (12) to be bounded.
The limitation of Lemma 3.5 is that it only applies to stationary policies. The lemma does not preclude the possibility that all stationary policies have unbounded returns but there exists a Markov policy with a bounded return. To show that this is impossible, we construct an upper bound on $\bm{w}^{t,\star}$ that decreases monotonically with $t$ and converges when bounded. The proof then concludes by squeezing $\bm{w}^{t,\star}$ between a lower bound that converges to the upper bound. We believe this technique allows us to relax the limiting assumptions from prior work (Patek 2001; de Freitas, Freire, and Delgado 2020).
Finally, note that our results imply that an optimal stationary policy exists whenever the planning horizon $T$ is sufficiently large. Because the set $\Pi_{\mathrm{MD}}$ is finite, one policy must be optimal for a sufficiently large $T$. This property mirrors turnpikes in discounted MDPs (Puterman 2005).
@@ 3.3 Algorithms
We now briefly describe the algorithms we use to compute the optimal ERM-TRC policies. Surprisingly, the major algorithms for discounted MDPs, including value iteration, policy iteration, and linear programming, can be adapted to this risk-averse setting with only minor modifications.
Value iteration is the most direct method for computing the optimal value function (Puterman 2005). The value iteration computes a sequence of $\bm{w}^{k},k=0,\dots$ such that
@@
The optimistic initialization of $\bm{w}^{0}$ is important and guarantees convergence directly from the monotonicity argument used to prove Theorem 3.3.
Policy iteration (PI) starts by initializing with a stationary policy $\pi_{0}=(\bm{d}^{0})_{\infty}\in\Pi_{\mathrm{SD}}$. Then, for each iteration $k=0,\dots$, the PI alternates between policy evaluation step and policy improvement step:
$\bm{w}^{k}$ $=-(\bm{I}-\bm{B}^{\bm{d}^{k}})\bm{b}^{\bm{d}^{k}},$
$\bm{d}^{k+1}$ $\in\operatorname*{argmax}_{\bm{d}\in\mathcal{D}}\bm{B}^{\bm{d}}\bm{w}^{k}-\bm{b}^{\bm{d}}.$
PI converges because it monotonically improves the value functions when initialized with a policy $\bm{d}^{0}$ with bounded return (Patek 2001). However, we are unaware of a practical approach to find such an initial policy.
@@
Here, $\bm{B}_{s,\cdot}^{a}=(\bm{B}_{s,s_{1}}^{a},\cdots,\bm{B}_{s,s_{S}}^{a})$, $\bm{B}_{s,s^{\prime}}^{a}$ and $b_{s}^{a}$ are constructed as in (11). We use the shorthand $\bm{B}^{a}=\bm{B}^{\bm{d}}$ and $\bm{b}^{a}=\bm{b}^{\bm{d}}$ where $d_{a^{\prime}}(s)=1$ if $a=a^{\prime}$ for each $s\in\mathcal{S},a^{\prime}\in\mathcal{A}$.
@@ 4 Solving EVaR Total Reward Criterion
This section shows that the EVaR-TRC objective can be reduced to a sequence of ERM-TRC problems, similarly to the discounted case (Hau, Petrik, and Ghavamzadeh 2023). As a result, an optimal stationary EVaR-TRC exists and can be computed using the methods described in Section 3.
@@
$\sup_{\pi\in\Pi_{\mathrm{HR}}}\liminf_{t\to\infty}\operatorname{EVaR}_{\alpha}^{\pi,\bm{\mu}}\left[\sum_{k=0}^{t}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right].$ (14)
In contrast with Ahmadi et al. (2021b), the objective in (14) optimizes the static rather than nested EVaR.
@@ 4.1 Reduction to ERM-TRC
$h_{t}(\pi,\beta)\;:=\;g_{t}(\beta)+\beta^{-1}\log(\alpha),$ (15)
@@
$\operatorname{EVaR}_{\alpha}^{\pi,\bm{\mu}}\left[\sum_{k=0}^{t}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]\;=\;\sup_{\beta>0}h_{t}(\pi,\beta),$ (16)
for each $\pi\in\Pi_{\mathrm{HR}}$ and $t\in\mathbb{N}$. However, note that the limit in the definition of $\sup_{\beta>0}h_{\infty}^{\star}(\pi,\beta)$ is inside the supremum unlike in the objective in (14).
@@
To circumvent the challenges with (16) mentioned above, we replace the supremum with a maximum over a finite set $\mathcal{B}(\beta_{0},\delta)$ of discretized $\beta$ values:
@@
$\beta_{k}\;:=\;\frac{\beta_{k}\log(\alpha)}{\beta_{k}\delta+\log(\alpha)},\quad\beta_{K}\;:=\;\frac{-\log(\alpha)}{\delta}.$ (17b)
The formulation in (17) resembles (19) and (20) in Hau, Petrik, and Ghavamzadeh (2023) but differs in the choice of $\beta_{0}$ because the TRC is unbounded with probability 1.
@@
where $\beta_{0}>0$ is chosen such that $g_{\infty}^{\star}(0)\leq g_{\infty}^{\star}(\beta_{0})-\delta$. Then the limits below exist and:
$\begin{gathered}\lim_{t\to\infty}\operatorname{EVaR}_{\alpha}^{\pi^{\star},\bm{\mu}}\left[\sum_{k=0}^{t}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]\\ \geq\sup_{\pi\in\Pi_{\mathrm{HR}}}\lim_{t\to\infty}\sup_{\beta>0}h(\pi,\beta)-\delta.\end{gathered}$ (18)
@@
There exists an optimal stationary policy $\pi^{\star}\in\Pi_{\mathrm{SD}}$ that attains the supremum in (14).

### v2

@@ 3 Solving ERM Total Reward Criterion
Our objective in this section is to maximize the ERM-TRC objective for some given $\beta>0$ defined as
$\sup_{\pi\in\Pi_{\mathrm{HR}}}\liminf_{t\to\infty}\operatorname{ERM}_{\beta}^{\pi,\bm{\mu}}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right].$ (5)
The definition employs limit inferior because the limit may not exist for non-stationary policies. Return functions $g_{t}\colon\Pi_{\mathrm{HR}}\times\mathbb{R}_{++}\to\mathbb{R}$ and $g_{t}^{\star}\colon\mathbb{R}_{++}\to\mathbb{R}$ for a horizon $t\in\mathbb{N}$ and the infinite-horizon versions $g_{t}\colon\Pi_{\mathrm{HR}}\times\mathbb{R}_{++}\to\bar{\mathbb{R}}$ and $g_{t}^{\star}\colon\mathbb{R}_{++}\to\bar{\mathbb{R}}$ are defined
$g_{t}(\pi,\beta)$ $:=\operatorname{ERM}_{\beta}^{\pi,\bm{\mu}}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right],$ (6)
$g_{t}^{\star}(\beta)$ $:=\sup_{\pi\in\Pi_{\mathrm{HR}}}g_{t}(\pi,\beta),$
$g_{\infty}(\pi,\beta)$ $:=\liminf_{t\to\infty}g_{t}(\pi,\beta),$
$g_{\infty}^{\star}(\beta)$ $:=\liminf_{t\to\infty}g_{t}^{\star}(\beta).$
Note that the functions $g_{\infty}$ and $g_{\infty}^{\star}$ can return infinite values and that (5) differs from $g^{\star}_{\infty}$ in the order of the limit and the supremum. Finally, when $\beta=0$, we assume that all $g$ functions are defined as the expectation. In the remainder of the section, we assume that the risk level $\beta>0$ is fixed and omit it in notations when its value is unambiguous from the context.
@@ 3.1 Finite Horizon
Finite-horizon functions $\bm{v}^{t}(\pi)\in\mathbb{R}^{S}$ and $\bm{v}^{t,\star}\in\mathbb{R}^{S}$ are defined for each horizon $t\in\mathbb{N}$ and policy $\pi\in\Pi_{\mathrm{MD}}$, $s\in\mathcal{S}$ as
@@
$v^{t,\star}_{s}$ $\;:=\;\max_{\pi\in\Pi_{\mathrm{MD}}}v^{t}_{s}(\pi),$
and $v_{e}^{t}(\pi):=0$.
Because the nonlinearity of ERM complicates the analysis, it will be convenient to instead rely on exponential value function $\bm{w}^{t}(\pi)\in\mathbb{R}^{S}$ for $\pi\in\Pi_{\mathrm{MD}}$, $t\in\mathbb{N}$, and $s\in\mathcal{S}$ that satisfy
@@
The optimal $\bm{w}^{t,\star}\in\mathbb{R}^{S}$ is defined analogously from $\bm{v}^{t,\star}$. Note that $\bm{w}^{t}<\bm{0}$ (componentwise) and $\bm{w}^{0}(\pi)=\bm{w}^{0,\star}=-\bm{1}$ for any $\pi\in\Pi_{\mathrm{MD}}$. Similar exponential value functions have been used previously in exponential utility function objectives Denardo and Rothblum (1979); Patek (2001), in the analysis of robust MDPs, and even in regularized MDPs (see Grand-Clément and Petrik (2022) and references therein).
@@
$L^{\star}\bm{w}$ $\;:=\;\max_{\bm{d}\in\mathcal{D}}L^{\bm{d}}\bm{w}=\max_{\bm{d}\in\operatorname{ext}\mathcal{D}}L^{\bm{d}}\bm{w},$
@@
For each $t=1,\dots$, and $\pi=(\bm{d}_{0},\dots,\bm{d}_{t-1})\in\Pi_{\mathrm{MR}}^{t}$, the exponential values satisfy that
@@
for some $\pi^{\star}\in\Pi_{\mathrm{MD}}^{t}$.
The proof of Theorem 3.1 is standard and has been established both in the context of ERMs Hau et al. (2023b) and utility functions Patek (1997).
The following corollary follows directly from Theorem 3.1 by algebraic manipulation and by the monotonicity of exponential value function transformation and the ERM.
@@ 3.2 Infinite Horizon
We now turn to construct infinite-horizon optimal policies as a limiting case of the finite horizon. An important quantity is the infinite-horizon exponential value function defined for each $\pi\in\Pi_{\mathrm{HR}}$ as
$\bm{w}^{\infty}(\pi)\;:=\;\liminf_{t\to\infty}\bm{w}^{t}(\pi),\quad\bm{w}^{\infty,\star}\;:=\;\liminf_{t\to\infty}\bm{w}^{t,\star}.$
Note again that we use the inferior limit because the limit may not be defined for non-stationary policies. The limiting infinite-horizon value functions $\bm{w}^{\infty}(\pi)$ and $\bm{w}^{\infty,\star}$ are defined analogously from $\bm{v}^{t}(\pi)$ and $\bm{v}^{t,\star}$ using the inferior limit. The following theorem is the main result of this section. It shows that for an infinite horizon, the optimal exponential value function is attained by a stationary deterministic policy and is a fixed point of the exponential Bellman operator.
@@
Asuming the hypothesis of Theorem 3.3, we have that $\bm{v}^{\infty,\star}=\bm{v}^{\infty}(\pi^{\star})$ and
@@
We now outline the proof of Theorem 3.3; see Su et al. (2024a, appendix D.4) for details. To establish Theorem 3.3, we show that $\bm{w}^{t,\star}$ converges to a fixed point as $t\to\infty$. Standard arguments do not apply to our setting Puterman (2005); Kallenberg (2021); Patek (2001) because the ERM-TRC Bellman operator is not an $L_{\infty}$-contraction, it is not linear, and the values in value iteration do not increase or decrease monotonically. Although the exponential Bellman operator $L^{\bm{d}}$ is linear, it may not be a contraction.
@@
The value $\bm{z}$ can be interpreted as the exponential value function at the termination of the process following $\pi$ for $t$ periods. Note that $\bm{w}^{t}(\pi)=\bm{w}^{t}(\pi,\bm{1}),\,\forall\pi\in\Pi_{\mathrm{MR}},t\in\mathbb{N}.$
An important technical result we show is that the only way a stationary policy’s return can be bounded is if the policy’s matrix has a spectral radius strictly less than $1$.
@@
Lemma 3.5 uses the transience property to show that the Perron vector (with the maximum absolute eigenvalue) $\bm{f}$ of $\bm{B}^{\bm{d}}$ satisfies that $\bm{f}^{\top}\bm{b}^{\bm{d}}>0$. Therefore, $\rho(\bm{B}^{\bm{d}})<1$ is necessary for the series in (12) to be bounded.
The limitation of Lemma 3.5 is that it only applies to stationary policies. The lemma does not preclude the possibility that all stationary policies have unbounded returns, but a Markov policy with a bounded return exists. We construct an upper bound on $\bm{w}^{t,\star}$ that decreases monotonically in $t$ and converges to show this is impossible. The proof then concludes by squeezing $\bm{w}^{t,\star}$ between a lower and the upper bound with the same limits. This technique allows us to relax the limiting assumptions from prior work Patek (2001); de Freitas et al. (2020). Finally, our results imply an optimal stationary policy exists whenever the planning horizon $T$ is sufficiently large. Because the set $\Pi_{\mathrm{SD}}$ is finite, one policy must be optimal for a sufficiently large $T$. This property suggests behavior similar to turnpikes in discounted MDPs Puterman (2005).
@@ 3.3 Algorithms
We now briefly describe the algorithms we use to compute the optimal ERM-TRC policies. Surprisingly, the main algorithms for discounted MDPs, including value iteration, policy iteration, and linear programming, can be adapted to this risk-averse setting with only minor modifications.
Value iteration is the most direct method for computing the optimal value function Puterman (2005). The value iteration computes a sequence of $\bm{w}^{k},k=0,\dots$ such that
@@
The initialization of $\bm{w}^{0}=\bm{0}$ is essential and guarantees convergence directly from the monotonicity argument used to prove Theorem 3.3.
Policy iteration (PI) starts by initializing with a stationary policy $\pi_{0}=(\bm{d}^{0})_{\infty}\in\Pi_{\mathrm{SD}}$. Then, for each iteration $k=0,\dots$, PI alternates between the policy evaluation step and the policy improvement step:
$\bm{w}^{k}=-(\bm{I}-\bm{B}^{\bm{d}^{k}})^{-1}\bm{b}^{\bm{d}^{k}},\;\bm{d}^{k+1}\in\operatorname*{argmax}_{\bm{d}\in\mathcal{D}}\bm{B}^{\bm{d}}\bm{w}^{k}-\bm{b}^{\bm{d}}.$
PI converges because it monotonically improves the value functions when initialized with a policy $\bm{d}^{0}$ with bounded return Patek (2001). However, we lack a practical approach to finding such an initial policy.
@@
Here, $\bm{B}_{s,\cdot}^{a}=(B_{s,s_{1}}^{a},\cdots,B_{s,s_{S}}^{a})$, $B_{s,s^{\prime}}^{a}$ and $b_{s}^{a}$ are constructed as in (11). We use the shorthand $\bm{B}^{a}=\bm{B}^{\bm{d}}$ and $\bm{b}^{a}=\bm{b}^{\bm{d}}$ where $d_{a^{\prime}}(s)=1$ if $a=a^{\prime}$ for each $s\in\mathcal{S},a^{\prime}\in\mathcal{A}$.
It is important to note that the value functions, as well as the coefficients of $\bm{B}^{\bm{d}}$ may be irrational. It is, therefore, essential to study the sensitivity of the algorithms to errors in the input. However, this question is beyond the scope of the present paper, and we leave it for future work.
@@ 4 Solving EVaR Total Reward Criterion
This section shows that the EVaR-TRC objective can be reduced to a sequence of ERM-TRC problems, similarly to the discounted case Hau et al. (2023b). As a result, an optimal stationary EVaR-TRC policy exists and can be computed using the methods described in Section 3.
@@
$\sup_{\pi\in\Pi_{\mathrm{HR}}}\liminf_{t\to\infty}\operatorname{EVaR}_{\alpha}^{\pi,\bm{\mu}}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right].$ (14)
In contrast with Ahmadi et al. (2021b), the objective in (14) optimizes EVaR rather than Nested EVaR.
@@ 4.1 Reduction to ERM-TRC
$h_{t}(\pi,\beta)\;:=\;g_{t}(\pi,\beta)+\beta^{-1}\log(\alpha),$ (15)
@@
$\operatorname{EVaR}_{\alpha}^{\pi,\bm{\mu}}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]\;=\;\sup_{\beta>0}h_{t}(\pi,\beta),$ (16)
for each $\pi\in\Pi_{\mathrm{HR}}$ and $t\in\mathbb{N}$. However, note that the limit in the definition of $\sup_{\beta>0}h_{\infty}^{\star}(\beta)$ is inside the supremum unlike in the objective in (14).
@@
To circumvent the challenges described above, we replace the supremum in (16) with a maximum over a finite set $\mathcal{B}(\beta_{0},\delta)$ of discretized $\beta$ values:
@@
$\beta_{k+1}\;:=\;\frac{\beta_{k}\log\frac{1}{\alpha}}{\log\frac{1}{\alpha}-\beta_{k}\delta},\qquad\beta_{K}\;\geq\;\frac{\log\frac{1}{\alpha}}{\delta},$ (17b)
for an appropriately chosen value $K$ for each $\beta_{0}$ and $\delta$. We assume that the denominator in the expression for $\beta_{k+1}$ in Equation 17b is positive; otherwise $\beta_{k+1}=\infty$ and $\beta_{k}$ is sufficiently large.
The construction in (17) resembles equations (19) and (20) in Hau et al. (2023b) but differs in the choice of $\beta_{0}$ because Hoeffding’s lemma does not readily bound the TRC criterion.
The following proposition upper-bounds the value of $K$; see (Hau et al., 2023b, theorem 4.3) for a proof that $K$ is polynomial in $\delta$.
Proposition 4.1.
Assume a given $\beta_{0}>0$ and $\delta\in(0,1)$ such that $\beta_{0}\delta<\log\frac{1}{\alpha}$. Then, to satisfy the condition in (17b), it is sufficient to choose $K$ as
$K:=\frac{\log z}{\log(1-z)},\quad\text{where}\quad z:=\frac{\beta_{0}\delta}{\log\frac{1}{\alpha}}.$ (18)
@@
where $\beta_{0}>0$ is chosen such that $g_{\infty}^{\star}(0)\leq g_{\infty}^{\star}(\beta_{0})-\delta$. Then the limits below exist and satisfy:
$\lim_{t\to\infty}\operatorname{EVaR}_{\alpha}^{\pi^{\star},\bm{\mu}}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]\geq\sup_{\pi\in\Pi_{\mathrm{HR}}}\lim_{t\to\infty}\sup_{\beta>0}h(\pi,\beta)-\delta.$ (19)
@@
There exists an optimal stationary deterministic policy $\pi^{\star}\in\Pi_{\mathrm{SD}}$ that attains the supremum in (14).